a boat travels 132 miles in 12 hours(with a constant speed).how much time will it take traveling 286 miles
Answer by jim_thompson5910(35256)

d = rt

132 = r*12

132/12 = r

11 = r

r = 11

So the constant speed is 11 mph

d = rt

286 = 11t

286/11 = t

26 = t

t = 26

So it will take 26 hours to go 286 miles.